OpenStudy (anonymous):

how do i write an equation of a line throught the given points: (-3,-2) and (0,-6)

OpenStudy (anonymous):

let (-3,-2) = (x1, y1) and let (0,-6) = (x2, y2)... slope of your line = m = (y2-y1)/(x2-x1) = -4/3... using point-slope form.... y-y2 = m(x-x2)... y-(-6) = (-4/3)(x-0)...y+6 = (-4/3)x...y = (-4/3)x-6....

OpenStudy (anonymous):

yes but how do i find the y intercept? where is that in the problem?

OpenStudy (anonymous):

y = mx + b...so y-int = b...so y = (-4/3)x -6...y-int = -6...or just let x = 0...and y = -6...